OpenStudy (anonymous):

Monochromatic light of 605 nm falls on a single slit, which is located 85 cm from a screen. The first dark band appears at 5.4 mm from the center. How wide is the slit?

OpenStudy (anonymous):

This is case of fraunhoffer single slit diffraction. calculate \[\sin \theta = 5.4/\sqrt{85^{2}+0.54^{2}}\] \[\beta=\pi*b*\sin \theta/\] condition for first maxima is \[\tan \beta=\beta\] which occurs at \[\beta=1.43*\pi\]